OpenStudy (cathyangs):

What is the intersection of the solution sets of the inequalities (2x-1)/(x+3)>_4 and (x+1)/(x+3)<_5 (>_is greater than or equal to) the answer choices all have different types of parentheses...

OpenStudy (anonymous):

Hmm...After reading this question again I am unsure if you have written it correctly. Please check it again :)

myininaya (myininaya):

Solve both inequalities. [(2x-1)-4(x+3)]/(x+3)>=0 and [(x+1)-5(x+3)]/(x+3)<=0 after getting both solutions sets to both problems find the intersection of those sets

myininaya (myininaya):

\[\frac{-(2x+13)}{x+3} \ge 0 \text{ and } -2 \cdot \frac{2x+7}{x+3} \le 0\]

myininaya (myininaya):

by the way x cannot be -3

OpenStudy (cathyangs):

It's correct...but I got that there is no intersection. the answer choices are: (-6.5, -3) [-6.5,-3.5] (-6.5,-3.5) [-6.5,-3.5) [-6.5,-3)

myininaya (myininaya):

so doing the first inequality we need to proceed as follows: \[\frac{-(2x+13)}{x+3} \ge 0\] so undefined at x=-3 and zero at x=-6.5 ---|---|--- -6.5 -3 now we choose a number in each of the intervals above

myininaya (myininaya):

if you want we can call \[f(x)=\frac{-(2x+13)}{x+3}\] so we can choose numbers to test like x=-8,-4,0 (you didn't have to choose these numbers-just find numbers in those three intervals)

myininaya (myininaya):

\[f(-8)=- ; f(-4)=+ ; f(0)=-\]

myininaya (myininaya):

we wanted to find when f>=0

myininaya (myininaya):

remember f(-4) was the only one to give us positive output

myininaya (myininaya):

so the interval [-6.5,-3) is the solution to first inequality now what did you get for the second inequality?

myininaya (myininaya):

or how about you try the second inequality?

OpenStudy (cathyangs):

I got -6.5 and -3 as the interval for the two...but how do you read the parentheses?

myininaya (myininaya):

[ ] these brackets mean include the endpoint ( ) these paraethesis mean exclude the endpoint so remember we didn't want to include -3 but we do want to include -6.5 so the answer is [-6.5,-3)

myininaya (myininaya):

remember we wanted to find when f>=0 f=0 when x=-6.5 f is undefined at x=-3

OpenStudy (cathyangs):

ah... thank you so much! :D

myininaya (myininaya):

ok but we still need to solve the second inequality

OpenStudy (cathyangs):

... doesn't it have the same solution? i am confuzzled. :\

myininaya (myininaya):

\[\text{ Let } g(x)=-2 \cdot \frac{2x+7}{x+3}\] we wanted to find g(x)<=0 so we see that g is undefined at x=-3 and we see that g is zero at x=-3.5

myininaya (myininaya):

no it doesn't have the same solution set

myininaya (myininaya):

It has a different zero is one reason we have different solution sets.

myininaya (myininaya):

it also a totally different function.

myininaya (myininaya):

So anyways we still have 3 intervals to test since we have 1 undefined place in the function and 1 zero in the function ---|----|---- -3.5 -3 so I would choose numbers and you don't have to choose the same numbers as me just numbers in the intervals above like x=-5,-3.2,0

OpenStudy (cathyangs):

... i graphed the two on a cartesian plane... the solution I got was [-6.5.-3) ?? very confuzzled.

myininaya (myininaya):

so remember \[\text{ Let } g(x)=-2 \cdot \frac{2x+7}{x+3} \] \[g(-5)=- ;g(-3.2)=+ ; g(0)=-\] so we are looking for when g<=0 and we see that we have - solutions for the following intervals (which makes the following intervals the solution set to the inequality) \[(-\infty,-3.5] \cup (-3,\infty)\]

myininaya (myininaya):

So solution from function f>=0 was \[[-6.5,-3)\] And solution from g<=0 was \[(-\infty,-3.5] \cup (-3,\infty) \] I think I will draw a rough graph so you can clearly see the intersection -----|------|-----|------ -6.5 -3.5 -3 *~~~~~~~~~() <----f>=0 ~~~~~~~~~~* ()~~~~~ <----g<=0 so intersection is \[[-6.5,-3.5]\]

OpenStudy (cathyangs):

thank you! :D

myininaya (myininaya):

hey one more thing...

myininaya (myininaya):

so in both of those inequalities there is equals part right?

OpenStudy (cathyangs):

it's blank than or equal to for both signs. sorry for not answering earlier.

myininaya (myininaya):

what?

myininaya (myininaya):

does _ this mean =?

myininaya (myininaya):

I'm going to assume <_ meant <= and >_ means >=